OpenStudy (anonymous):

show that if y1 and y2 be the solutions of the equation dy/dx+Py=Q, where P and Q are functions of x alone, and y2=y1z , then z= 1+ ae^(-Q/y1)dx.a being arbitrary constant

OpenStudy (usukidoll):

differential equation....

OpenStudy (anonymous):

yes

OpenStudy (usukidoll):

aww man I'm bad at proofs and I just went through a lot of math...umm you could ask phi, KingGeorge, OOOPS...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

z=1+ae^{(-\frac{Q}{y_1})}dx ?

OpenStudy (anonymous):

\[z=1+ae^{(-\frac{Q}{y_1})}dx \]?

OpenStudy (anonymous):

yes we have to prove it

OpenStudy (anonymous):

can you help me in this?

OpenStudy (anonymous):

I can get most of the way there, but the ending isn't working great... I'll show ya wht I have

OpenStudy (anonymous):

yes pleaase

OpenStudy (anonymous):

\[ \frac{\mathrm{d}y}{\mathrm{d}x}+Py = Q\] y1 and y2 are solutions, so \[ \frac{\mathrm{d}y_1}{\mathrm{d}x}+Py_1 = Q \quad \quad ; \quad P= \Big (Q-\frac{\mathrm{d}y_1}{\mathrm{d}x}\Big) \frac{1}{y_1}\] and \[ \frac{\mathrm{d}y_2}{\mathrm{d}x}+Py_2 = Q \quad \quad ; \quad P= \Big(Q-\frac{\mathrm{d}y_2}{\mathrm{d}x}\Big) \frac{1}{y_2}\] so we can equate the Ps \[\Big(Q-\frac{\mathrm{d}y_1}{\mathrm{d}x}\Big) \frac{1}{y_1}= \Big(Q-\frac{\mathrm{d}y_2}{\mathrm{d}x}\Big) \frac{1}{y_2}\] using y2 = zy1 gives \[\Big(Q-\frac{\mathrm{d}y_1}{\mathrm{d}x}\Big) \frac{1}{y_1}= \Big(Q-\frac{\mathrm{d}zy_1}{\mathrm{d}x}\Big) \frac{1}{zy_1}\] the y1s cancel out, bring the z to the other side, and you can use the product rule on d zy1/dx \[z\Big(Q-\frac{\mathrm{d}y_1}{\mathrm{d}x}\Big)= \Big(Q- \big(y_1\frac{\mathrm{d}z}{\mathrm{d}x}+z\frac{\mathrm{d}y_1}{\mathrm{d}x}\big)\Big)\] \[zQ-z\frac{\mathrm{d}y_1}{\mathrm{d}x}= Q- y_1\frac{\mathrm{d}z}{\mathrm{d}x}-z\frac{\mathrm{d}y_1}{\mathrm{d}x}\] the zdy1/dx's cancel out and you can move the Qs to one side \[zQ-Q= - y_1\frac{\mathrm{d}z}{\mathrm{d}x}\] \[-\frac{Q}{y_1}(z-1)= \frac{\mathrm{d}z}{\mathrm{d}x}\] \[-\frac{Q}{y_1}\mathrm{d}x= \frac{1}{(z-1)}\mathrm{d}z\] then you can integrate the right hand side to get ln(z-1), but I dunno how to keep the dx on the bottom on the left, or how to get the constant out front \[-\int \frac{Q}{y_1}\mathrm{d}x= \int \frac{1}{(z-1)}\mathrm{d}z\] \[-\int \frac{Q}{y_1}\mathrm{d}x=ln(z-1)+A\]

OpenStudy (anonymous):

any thoughts? :P

OpenStudy (anonymous):

let me see wait

OpenStudy (anonymous):

It's just sooo close I think it's the right way to go, just dunno where to go :P

OpenStudy (anonymous):

yes its right i m also stuck here

OpenStudy (anonymous):

but thanx ,,

OpenStudy (anonymous):

I mean you can get it to \[-\int \frac{Q}{y_1}\mathrm{d}x=ln(z-1)+A\] \[-\int \frac{Q}{y_1}\mathrm{d}x-A=ln(z-1)\] \[e^{-\int \frac{Q}{y_1}\mathrm{d}x-A}=z-1\] \[e^{-\int \frac{Q}{y_1}\mathrm{d}x}e^{-A}=z-1\] \[ \quad \quad e^{-A} = a\] \[z=1+ae^{-\int \frac{Q}{y_1}\mathrm{d}x}\] but there's still that dumb integral :P

OpenStudy (anonymous):

T_T

OpenStudy (anonymous):

yes may be there is mistake in the question

OpenStudy (anonymous):

I think that's the better answer ^_^

OpenStudy (anonymous):

becoz this is correct answer

OpenStudy (anonymous):

yes thanx a lot for your help

OpenStudy (anonymous):

welcome ^_^ That was a fun one :D

OpenStudy (anonymous):

:)